Find the number of ways in which seven different toys can be given to three children, if the youngest is to receive three toys and the others receive two toys each.
Answer by Edwin McCravy(20054)

Think of the 3 children lined up from youngest to oldest:

Choose 3 toys from the 7 to give the youngest in 7C3 ways.
Choose 2 of the remaining 4 toys to give to the next oldest in 4C2 ways.
Choose the remaining 2 toys to give to the oldest in 2C2 = 1 way.

Answer: 7C3*4C2*2C2 = 35*6*1 = 210 ways.

Edwin